### v2

# ACTIVE: Continuous Similarity Search for Vessel Trajectories

Abstract.
Publicly available vessel trajectory data is emitted continuously from the global AIS system. Continuous trajectory similarity search on this data has applications in, e.g., maritime navigation and safety. Existing proposals typically assume an offline setting and focus on finding similarities between complete trajectories. Such proposals are less effective when applied to online scenarios, where similarity comparisons must be performed continuously as new trajectory data arrives and trajectories evolve. We therefore propose a real-time continuous trajectory similarity search method for vessels (ACTIVE). We introduce a novel similarity measure, object-trajectory real-time distance, that emphasizes the anticipated future movement trends of vessels, enabling more predictive and forward-looking comparisons. Next, we propose a segment-based vessel trajectory index structure that organizes historical trajectories into smaller and manageable segments, facilitating accelerated similarity computations. Leveraging this index, we propose an efficient continuous similar trajectory search (CSTS) algorithm together with a variety of search space pruning strategies that reduce unnecessary computations during the continuous similarity search, thereby further improving efficiency. Extensive experiments on two large real-world AIS datasets offer evidence that ACTIVE is capable of outperforming state-of-the-art methods considerably. ACTIVE significantly reduces index construction costs and index size while achieving a 70% reduction in terms of query time and a 60% increase in terms of hit rate.
Tiantian Liu and Hengyu Liu have equal contributions.
Tianyi Li is the corresponding author.
1. Introduction
Automatic Identification System (AIS) data, which provides real-time information on vessel positions, speeds, and directions, offers a rich resource for analyzing vessel movements. With the growing availability of AIS data, understanding vessel trajectories has become essential for navigation safety (Li et al., 2024b) and environmental monitoring (Jin et al., 2021). Trajectory similarity search (Hu et al., 2023) plays a key role in applications such as collision avoidance (Li et al., 2022), route optimization (Xu et al., 2024), and movement pattern analysis (Hu et al., 2024).
Given the dynamic nature of maritime environments, real-time continuous trajectory similarity search is particularly valuable. It enables timely and proactive decision-making by identifying similar movement patterns as new AIS data arrives, facilitating responsive navigation and operational safety. Unlike traditional methods, which rely on retrospective analysis of complete trajectories, real-time similarity search operates incrementally, processing incoming data without requiring an entire trajectory. Moreover, real-time search considers the future movement trends and the destination of a moving object. For instance, as illustrated in Figure 1, for a moving vessel $o$, given its current trajectory $T_{o}$ and target destination $p_{o}^{d}$, we continuously search for similar historical trajectories, accounting for both past movements and the vessel’s future trend. Despite significant progress in real-time trajectory similarity search, existing studies (Ding et al., 2008; Xie et al., 2017; Wang et al., 2018; Li et al., 2020; He et al., 2022; Luo and Zeng, 2023; Bakalov et al., 2005b, a; Ta et al., 2017; Shang et al., 2017) face limitations, leading to the following challenges.
Figure 1. Example of trajectories at timestamp $t$
Challenge 1: How to effectively measure trajectory similarity. Existing methods, such as Dynamic Time Warping (Yi et al., 1998; Ying et al., 2016), Longest Common Subsequence (Vlachos et al., 2002), Fréchet distance (Alt and Godau, 1995; De Berg et al., 2013; Agarwal et al., 2014), Hausdorff distance (Hangouet, 1995; Bai et al., 2011), and Edit distance-based measures (Chen and Ng, 2004; Chen et al., 2005; Ranu et al., 2015), are commonly used to identify similar trajectories by analyzing historical movement data. However, they rely heavily on static or full-path comparisons, making them of low utility for evolving trajectories that need to be analyzed incrementally as new data arrives. Next, they focus exclusively on historical movements, overlooking the importance of an object’s near-future movement and target destination. This limits their effectiveness in real-time applications, such as collision detection, traffic management, and route prediction, where anticipating future movements is crucial. For example, in Figure 1, the most similar trajectory for $T_{o}$ is $T_{3}$ using Fréchet or Hausdorff distance. However, when considering the destination of the object, the most similar trajectory should be $T_{1}$.
Challenge 2: How to improve efficiency and scalability in continuous similarity trajectory search. Existing similar trajectory search methods (Ding et al., 2008; Xie et al., 2017; Wang et al., 2018; Li et al., 2020; He et al., 2022; Luo and Zeng, 2023) are usually designed for searches using traditional similarity metrics intended for historical data, making them less effective for real-time applications. Moreover, they are designed for offline, batch processing on static datasets, requiring full re-computations when new data arrives. This is unsuitable in a setting with continuous updates to AIS data.
Achieving efficient, real-time continuous similarity searches for vessel trajectories involves addressing several interconnected challenges. First, vessel trajectories are complex, often involving irregular and dynamic movements in open water. In contrast, on road networks, structured paths and intersections simplify trajectory analysis. Second, pruning irrelevant or low-similarity candidates with high accuracy is particularly complex in a dynamic environment where trajectory data changes continuously. Effective pruning methods must strike a balance between reducing candidate sets and maintaining the accuracy of similarity computations. Third, scaling a system and retaining fast response times when handling continuous queries over extended periods adds another layer of difficulty. The system must be capable of managing large-scale, real-time data streams while maintaining high accuracy and adapting to varying query loads. Together, these challenges call for new innovations to achieve a scalable and accurate trajectory search system.
Contributions. In response to these challenges, we present a novel framework for real-time continuous trajectory similarity search for vessels. To address Challenge 1, we propose a novel similarity measure called object-trajectory real-time distance, which prioritizes the future trends of vessel movements and their target destinations while accounting also for past trajectory parts and the significance of individual trajectory points. This measure enables more accurate predictions of future vessel positions and behaviors, making it better suited for real-time decision-making in maritime navigation and traffic management.
To address Challenge 2, we propose a novel index structure called the segment-based vessel trajectory index, which organizes historical vessel trajectories by partitioning them into smaller, manageable segments using a Minimum Bounding Rectangle (MBR)-optimization method. This partitioning accelerates similarity computations by enabling a focus on the most relevant trajectory candidates, avoiding the need for exhaustive analyses of entire trajectories. Moreover, we design an efficient continuous similar trajectory search algorithm with four speed-up strategies that exploit geometric properties, utilize intermediate data, such as similarities computed at previous timestamps, and adjust trajectory granularity. These strategies efficiently eliminate irrelevant trajectory comparisons early in the search process while maintaining accuracy. Together, they optimize continuous searches and reduce computational costs notably. The contributions are summarized as follows.
  • We propose a framework for real-time continuous trajectory similarity search for vessels (ACTIVE). To the best of our knowledge, it is the first to support real-time similar trajectory search while considering the future trends of moving vessels.
  • We develop a new index structure, the segment-based vessel trajectory index (SVTI), which partitions and organizes historical vessel trajectories to accelerate similarity searches.
  • We introduce a novel similarity measure, the object-trajectory real-time distance (OTRD), which considers future vessel trajectory trends to enhance similarity accuracy.
  • We present an efficient continuous similar trajectory search (CSTS) algorithm, that integrates strategies to reduce unnecessary computations and optimize the search process.
  • Experiments on two large real-world AIS datasets offer evidence that ACTIVE is capable of outperforming state-of-the-art methods by up to 70% in terms of query time, 60% in terms of hit rate, and reduces index construction costs significantly.
The rest of the paper is organized as follows. Section 2 presents the problem and solution overview. Section 3 covers the segment-based vessel trajectory index. Section 4 details the new similarity measure. Section 5 presents the algorithms and speed-up strategies for continuous similar vessel trajectory search. Section 6 reports the experimental results. Section 7 reviews related works. Section 8 concludes the paper and offers future directions.
2. Preliminaries
Section 2.1 presents key definitions. Section 2.2 defines the research problem. Section 2.3 gives an overview of the ACTIVE framework. Frequently used notation is listed in Table 1.
Table 1. Notation
  Symbol Description
$T_{i}$ A trajectory
$s^{u}_{i}$ The $u^{\textit{th}}$ segment of trajectory $T_{i}$
$\mathcal{T}$ A set of trajectories
$T_{i}=\langle s^{1}_{i},...,s^{u}_{i}\rangle$ A list trajectory segments
$o$ A moving object
$\mathcal{D}(o,T_{i})$ The distance between a moving object $o$ and a trajectory $T_{i}$
$T_{i}^{*}$ The trajectory $T_{i}$ with the last element removed
$d(p^{j^{\prime}},p^{j})$ The Euclidean distance between points $p^{j^{\prime}}$ and $p^{j}$
$L(p^{j^{\prime}},p^{j})$ The line segment between points $p^{j^{\prime}}$ and $p^{j}$
2.1. Trajectory
Definition 0 (GPS Point).
A GPS point $p=(x,y,t)$ includes a location with longitude $x$ and latitude $y$ and timestamp $t$.
For simplicity, we refer to a GPS point as a point $(x,y)$ when it is in a snapshot at a timestamp.
Figure 2. ACTIVE framework
Definition 0 (Historical Trajectory).
A historical trajectory $T_{i}$ is a finite, time-ordered sequence $T_{i}=\langle p_{i}^{1},p_{i}^{2},...,p_{i}^{n}\rangle$. Next, $\mathcal{T}=\{T_{i}\mid 1\leq i\leq|\mathcal{T}|\}$ is a set of $|\mathcal{T}|$ historical trajectories.
In Figure 1, $\mathcal{T}=\{T_{1},T_{2},T_{3},T_{4},T_{5}\}$. We simplify $T_{i}$ to $T$ when the index $i$ is not relevant.
Definition 0 (Segment).
A segment $s_{i}=\langle p_{i}^{j^{\prime}},\dots,p_{i}^{j}\rangle$ of a trajectory $T_{i}$ is a continuous sub-sequence of $T_{i}$ defined by an index range $j^{\prime},j]$, where $1\leq j^{\prime}\leq j\leq n$.
Building on Definition 2.3, a trajectory $T_{i}$ can be represented as a sequence of segments $T_{i}=\langle s_{i}^{1},s_{i}^{2},\dots,s_{i}^{u}\rangle$, where $s_{i}^{u}$ is the $u^{\textit{th}}$ segment of $T_{i}$. For example, in Figure 1, trajectory $T_{1}$ consists of two segments, $s_{1}^{1}$ and $s_{1}^{2}$, thus $T_{1}=\langle s_{1}^{1},s_{1}^{2}\rangle$, where $s_{1}^{1}=\langle p_{1}^{1},p_{1}^{2},p_{1}^{3},p_{1}^{4}\rangle$ and $s_{1}^{2}=\langle p_{1}^{5},p_{1}^{6},p_{1}^{7},p_{1}^{8}\rangle$.
Definition 0 (Moving Object).
A moving object $o$ changes its position over time in a given space. The destination of an object $o$ is given by $p_{o}^{d}$. Next, $o_{t}=(p_{o}^{c},T_{o})$ represents $o$’s information at time $t$, where $p_{o}^{c}$ is the position of $o$ at time $t$ and $T_{o}=\langle p_{o}^{1},p_{o}^{2},\dots,p_{o}^{c}\rangle$ is the trajectory of $o$ up to time $t$.
Example 0.
In Figure 1, at time $t$, $o_{t}=(p_{o}^{6},\langle p_{o}^{1},p_{o}^{2},p_{o}^{3},p_{o}^{4},$ $p_{o}^{5},p_{o}^{6}\rangle)$, which means its current position is $p_{o}^{6}$ and its trajectory up to $t$ is $\langle p_{o}^{1},p_{o}^{2},p_{o}^{3},p_{o}^{4},$ $p_{o}^{5},p_{o}^{6}\rangle$.
Definition 0 (Pivotal Point).
Given a historical trajectory $T=\langle p^{1},p^{2},\dots,p^{n}\rangle$ and a moving object $o$ with $o_{t}=(p_{o}^{c},T_{o})$, the pivotal point $p^{x}$ of $T$ w.r.t. $o$ is the point in $T$ closest to $p_{o}^{c}$ at time $t$. Letting $p^{x}=p^{j}$, the original trajectory $T$ is represented as $(T^{\prime},T^{L})$, where $T^{\prime}=\langle p^{1},p^{2},\dots,p^{j}\rangle$ and $T^{L}=L(p^{x},p^{n})$ is the line segment between pivotal point $p^{x}$ and the last point $p^{n}$ in $T$.
Example 0.
Continuing Example 2.5, the pivotal point $p_{4}^{x}$ of $T_{4}$ w.r.t. $o$ is $p_{4}^{5}$, as it is the closest to $p_{o}^{6}$. Consequently, we use $(T_{4}^{\prime},T_{4}^{L})$ to represent $T_{4}$, where $T_{4}^{\prime}=\langle p_{4}^{1},p_{4}^{2},p_{4}^{3},p_{4}^{4},p_{4}^{5}\rangle$ and $T_{4}^{L}=L(p_{4}^{5},p_{4}^{7})$ is the dotted line between $p_{4}^{5}$ and $p_{4}^{7}$.
Definition 0 (Range Query).
Given a range $r$, a moving object $o$ with $o_{t}=(p_{o}^{c},T_{o})$, a range query $\mathcal{R}(p_{o}^{c},r)$ returns the set of historical trajectories $\mathcal{T}_{r}=\{T_{i}\mid T_{i}\in\mathcal{T}\wedge d(p_{o}^{c},p_{i}^{x})\leq r\}$, where $p_{i}^{x}$ is the pivotal point of trajectory $T_{i}$, and $d(p_{o}^{c},p_{i}^{x})$ is the distance between the current position $p_{o}^{c}$ of $o$ and the pivotal point $p_{i}^{x}$ of $T_{i}$.
Example 0.
Continuing Example 2.7, given the range $r$ in the example, the range query $\mathcal{R}(p_{o}^{c},r)$ returns $\mathcal{T}_{r}=\{T_{1},T_{3},T_{4}\}$.
2.2. Problem Definition
Research Problem 0 (Continuous Similar Trajectory Search (CSTS)).
[…]
Here, $\mathcal{D}(o,T)$ is the similarity measure OTRD (defined in Section 4.2). Continuing Example 2.9, given $k=2$, $\Theta_{t}=\{T_{1},T_{4}\}$.
2.3. Framework Overview
Figure 2 shows the three key modules in ACTIVE framework:
Segment-based Vessel Trajectory Index module. Historical trajectories are partitioned into smaller segments using an MBR-optimization method and organized in an index. This index captures connection information, enabling efficient reconstruction of full trajectories when needed and supports spatially optimized search.
Similarity Computation module. It computes the real-time distance between a moving object and historical trajectories using the novel similarity measure OTRD, which consists of the Historical Trajectory Distance (HTD) and the Target-Trajectory Distance (TTD). The OTRD measure prioritizes the future trends of vessel movements and the target destination while also considering the relevance of trajectory segments and the weighted significance of individual points.
Figure 3. An example of SVTI
Continuous Similar Trajectory Search module. Using the proposed CSTS algorithm, this module performs continuous search with four speed-up strategies that exploit geometric properties, utilize intermediate data, such as similarities computed at previous timestamps, and adjust trajectory granularity to reduce computational cost while preserving accuracy. Given a moving object $o$, a set of historical trajectories $\mathcal{T}$, a search range $r$, and an integer $k$, it continuously identifies the $k$ most similar trajectories $\Theta_{t}$ for $o$ at each timestamp $t$.
3. Segment-based Trajectory Index
3.1. Index Structure
The segment-based vessel trajectory index (SVTI) is a spatial access method designed to index vessel trajectory data efficiently. Each trajectory is divided into segments and a R-Tree structure organizes these segments hierarchically, enabling fast querying by grouping spatially close segments. Each leaf node stores the Minimum Bounding Rectangle (MBR) of its segments and a set of tuples, such as ($i,U$), where $i$ is the trajectory ID, and $U$ is the set of segment IDs within $T_{i}$. The leaf nodes also link to detailed information, including the sequence of points forming each segment and the search path for preceding segments, enabling seamless segment-to-trajectory mapping. Non-leaf nodes store the MBR of the node itself and references to trajectory and segment IDs in their child nodes. This hierarchical design preserves spatial relationships and trajectory continuity across different tree levels, ensuring efficient indexing and querying.
Example 0.
Figure 3 illustrates an example of SVTI. The index includes two leaf nodes, $M_{11}$ and $M_{12}$. In $M_{11}$, four segments are stored, each corresponding to one of four trajectories: $T_{1}$, $T_{2}$, $T_{3}$, $T_{4}$. $M_{12}$ contains three segments, involving trajectories $T_{1}$, $T_{3}$, $T_{4}$. Each leaf node is linked to both the segment data and the connection information. For example, $M_{12}$ contains $s^{2}_{1}$, which is the second segment in $T_{1}$. Detailed segment information is associated with $s^{2}_{1}$, including its preceding segment $s^{1}_{1}$ and the path $\langle M_{0},M_{11}\rangle$ that leads to $s^{1}_{1}$. The two leaf nodes, $M_{11}$ and $M_{12}$, are then grouped into a higher-level node $M_{0}$, which serves as the root node in this case. $M_{0}$ manages a total of seven segments, involving four different trajectories.
3.2. Segment Partitioning
The performance of the index is significantly influenced by both the number of segments and the total area covered by the MBRs. Reducing the total MBR area minimizes overlap between nodes, thereby improving the efficiency of spatial indexing and query performance. Therefore, the goal of segment partitioning is to split a trajectory such that the cumulative area of the MBRs for the resulting segments is minimized.
The segment partitioning method uses dynamic programming to determine the optimal segmentation. For each point $p^{j}$ in a given trajectory, the minimum total MBR area, $F[j]$, is computed by considering all possible previous points, starting from $p^{j-l+1}$, where $l_{\textit{min}}\leq l\leq l_{\textit{max}}$. Here, $l_{\textit{min}}$ and $l_{\textit{max}}$ denote the minimum and maximum segment lengths, respectively. The recurrence relation is defined as follows:
(1) $F[j]=\textit{min}\{F[j-l]+\mathcal{M}_{\textit{area}}(p^{j-l+1},p^{j})|l_{\textit{min}}\leq l\leq l_{\textit{max}}\},$
where $\mathcal{M}_{\textit{area}}(p^{j-l+1},p^{j})$ is the MBR area of segment $\langle p^{j-l+1},...,p^{j}\rangle$. Parameter $l_{\textit{min}}$ plays a critical role in controlling the number of segments, which directly impacts the construction of the R-tree-based index structure. By reducing the number of segments, the efficiency of tree indexing and query processing is improved. Meanwhile, $l_{\textit{max}}$ defines an upper bound on the time complexity of pruning operations. This ensures that the computational cost remains manageable, even as the trajectory data becomes more complex. By balancing segment length constraints and MBR optimization, the partitioning process enhances spatial query performance and overall system efficiency in managing large-scale trajectory data.
4. Similarity Measure
4.1. Case Analysis
Existing trajectory distance measures, such as Dynamic Time Warping (DTW), Hausdorff distance, and Fréchet distance, typically evaluate the entire trajectory and treat each point with equal significance. However, these methods exhibit certain limitations when applied to real-time trajectory distance computations. Consider the following cases:
Figure 4. Examples of gaps between existing similarity measures and real needs
Case 1: Trend Neglect. Traditional trajectory similarity measures focus on minimizing geometric or temporal differences without considering the directional movement. In Figure 4(a), the distance between trajectories $T_{1}$ and $T_{o}$ is computed to be the same as that between $T_{2}$ and $T_{o}$ under existing measures. However, in many real-world applications, such as predictive modeling or motion forecasting, the trajectory trend plays a critical role in determining similarity. For example, although $T_{2}$ is farther from $T_{o}$ at earlier timestamps, its trend more closely aligns with $T_{o}$’s current movement direction, suggesting that $T_{2}$ is more likely to converge with $T_{o}$ in the near future. This indicates that ignoring the trend can lead to suboptimal or misleading results, particularly in scenarios where the future behavior of trajectories is of primary interest.
Case 2: Segment Relevance. Conventional trajectory similarity measures assume that all segments or points along a trajectory contribute equally to the overall similarity computation. This assumption overlooks the fact that, in many downstream tasks, certain segments of a trajectory may carry more relevance than others. For example, in Figure 4(b), given a trajectory $T_{o}$ and a set of historical trajectories $T_{1},T_{2}$}, only the segment highlighted in the red box is of primary interest for comparison. Ignoring the varying importance of trajectory segments can lead to suboptimal decision-making in critical scenarios.
Case 3: Target Point Neglect. In many practical applications, such as logistics and real-time navigation, the target destination of a moving object is of paramount importance. Current trajectory similarity measures, however, are predominantly backward-looking, focusing on historical data while disregarding the significance of the target destination and the trajectory’s future course. For instance, as shown in Figure 4(c), while the historical paths of trajectories $T_{1}$ and $T_{2}$ are similar, $T_{1}$’ trend after the pivotal point of are more aligned with the projected path of $T_{o}$ when the target destination is taken into consideration. Incorporating the target point into trajectory similarity measures would not only enhance the accuracy of real-time trajectory search but also enable more reliable predictions of future movements in applications such as autonomous navigation, shipping logistics, and traffic management.
4.2. Object-Trajectory Real-time Distance
The object-trajectory real-time distance (OTRD) is a similarity measure function $\mathcal{D}(o,T_{i})$ to measure the distance between a moving object $o$ at timestamp $t$ with information $(p_{o}^{c},T_{o})$ and a trajectory $T_{i}$. The distance between $o$ and $T_{i}$ consists of the Historical Trajectory Distance (HTD) and the Target-trajectory Distance (TTD). Specifically, the HTD is the distance between $T_{o}$ of $o$ and $T^{\prime}_{i}$ of $T_{i}$, denoted as $\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$, where $T^{\prime}_{i}=\langle p_{i}^{1},\dots,p_{i}^{x}\rangle$. The TTD is the distance from $o$’ target destination $p_{o}^{d}$ to the line $T^{L}=L(p_{i}^{x},p_{i}^{n})$, denoted as $\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L})$. The OTRD considers more about the points near the current position of the moving object. Besides, the target destination is also considered to improve the accuracy of the similarity in terms of the future trend.
Definition 0 (Object-Trajectory Real-time Distance).
Given a moving object $o$ at timestamp $t$ with information $(p_{o}^{c},T_{o})$ and a trajectory $T_{i}=\langle p_{i}^{1},p_{i}^{2},\dots,p_{i}^{n}\rangle$, the Object-Trajectory Real-time Distance between $o$ and $T_{i}$ is defined as follows:
(2) $\mathcal{D}(o,T_{i})=\alpha\cdot\mathcal{D}_{H}(T_{o},T^{\prime}_{i})+(1-\alpha)\cdot\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L})$
[…]
(4) $\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L})=d(p_{o}^{d},L(p_{i}^{x},p_{i}^{n}))$
Here, $d(p_{o}^{*},p_{i}^{*})$ is the euclidean distance between $p_{o}^{*}$ and $p_{i}^{*}$, where $p_{o}^{*}$ and $p_{i}^{*}$ are last points in $T_{o}$ and $T^{\prime}_{i}$. $\theta\in[0,1]$ is the decay factor. $x=\text{argmin}_{0<j\leq n}d(p_{o}^{c},p_{i}^{j})$ is the ID of the pivotal point in $T_{i}$. $T_{o}^{*}$ and $T^{\prime}_{i}$$^{*}$ are the segments of $T_{o}$ and $T^{\prime}_{i}$ with the last point removed. The distance can be flexibly customized by a trade-off parameter $\alpha\in[0,1]$ according to specific application needs. The process of OTRD computation is detailed in Section 5.2.
Figure 5. Example of OTRD measure
Example 0.
Figure 5 shows an example of OTRD. Given a trajectory $T_{1}=\langle(1,2),(3,3),(5,3),(6,5),(7,6),(9,7)\rangle$, a moving object $o$ at time $t$ with the target destination $p_{o}^{d}=(9,5)$ and $(p_{o}^{c},T_{o})$, where $p_{o}^{c}=(6,3)$ and $T_{o}=\langle(2,1),(4,2),(5,2),(6,3)\rangle$, the trade-off parameter $\alpha=0.9$, and the decay factor $\theta=0.9$, we have the pivotal point of $T_{1}$, $p_{1}^{x}=p_{1}^{3}=(5,3)$, and two parts of $T_{1}$, $T^{\prime}_{1}=\langle(1,2),(3,3),(5,3)\rangle$ and $T_{1}^{L}=L((5,3),(9,7))$. The HTD and TTD can be computed accordingly, i.e., $\mathcal{D}_{H}(T_{o},T^{\prime}_{1})=1.14$ and $\mathcal{D}_{T}(p_{o}^{d},T_{1}^{L})=1.41$. Therefore, the OTRD between $o$ and $T_{1}$ is $\mathcal{D}(o,T_{1})=0.9*1.14+0.1*1.41=1.167$.
5. Search Algorithm
5.1. Speed-up Strategies
To enhance the efficiency of similarity computation and continuous similar trajectory search, we propose four speed-up strategies. These strategies leverage geometric properties, intermediate data, and trajectory granularity adjustments to reduce computational overhead while preserving accuracy.
The similarity computation operates segmented trajectories to improve the efficiency. The MBR of a segment $s$ denoted as $\mathcal{M}(s)=(v_{l}^{u},v_{l}^{b},v_{r}^{u},v_{r}^{b})$, where $v_{l}^{u}$ and $v_{l}^{d}$ are the upper-left and the bottom-left vertices of the MBR, while $v_{r}^{u}$ and $v_{r}^{d}$ are the upper-right and the bottom-right vertices. For a point $p$ outside of $\mathcal{M}(s)$, the distance between $p$ and $\mathcal{M}(s)$ is computed as follows:
(5) $\mathcal{D}_{s}(p,\mathcal{M}(s))=\text{min}\{d(p,v_{l}^{u}),d(p,v_{l}^{b}),d(p,v_{r}^{u}),d(p,v_{r}^{b})\}.$
Strategy 1 (Segment Pruning).
When computing the HTD between the moving object $o$ at timestamp $t$ with information $o_{t}=(p_{o}^{c},T_{o})$ and a trajectory $T_{i}$, a segment $s_{i}$ can be pruned if
(6) $\theta^{c-j}\mathcal{D}_{s}(p_{o}^{j},\mathcal{M}(s_{i}))>\textit{dist}_{\textit{min}},$
where $c$ is the ID of $o$’s the current point at timestamp $t$, $p_{o}^{j}$ is a point in $o$’s historical trajectory $T_{o}$, $T^{\prime}_{i}$ is the sub-trajectory of $T_{i}$ from the starting point to the pivotal point, $s_{i}$ is a segment in $T^{\prime}_{i}$, and $\textit{dist}_{\textit{min}}$ is the current minimum distance between $p_{o}^{j}$ and $T^{\prime}_{i}$.
Lemma 5.1.
Given a segment $s$ and a point $p$, the distance between $p$ and any point $p^{j}$ in $s$ is no less than $\mathcal{D}_{s}(p,\mathcal{M}(s))$.
Proof.
If $p^{j}$ lies inside the rectangle, the shortest possible path from $p$ to $p^{j}$ must pass through a boundary point because the shortest distance from an external point to a shape is to its boundary. This boundary point must be closer or equal to the nearest vertex $v_{\textit{min}}$. If $p^{j}$ lies on an edge, it lies between two vertices of the rectangle. Since the distance between $p$ and any point on the edge is bounded by the distance from $p$ to the closest vertex, the distance $d(p,p^{j})\geq d(p,v_{\textit{min}})=\mathcal{D}_{s}(p,\mathcal{M}(s))$. ∎
Strategy 2 ($k$-bound Pruning).
Given the current $k$-th smallest Object-Trajectory Real-time Distance $\textit{dist}_{k}$ between a moving object $o$ at timestamp $t$ with information $(p_{o}^{c},T_{o})$ and the computed historical trajectories, a trajectory $T_{i}$ can be pruned if
(7) $\mathcal{D}^{l}(p_{o}^{j},T^{\prime}_{i})\geq\frac{\textit{dist}_{k}-(1-\alpha)\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L})}{\alpha\cdot\theta^{c-j}},$
where
(8) $\mathcal{D}^{l}(p_{o}^{j},T^{\prime}_{i})=\text{min}\{d(p_{o}^{j},p_{i})|p_{i}\in T^{\prime}_{i}\}.$
Lemma 5.2.
Given a moving object $o$ at timestamp $t$ with information $o_{t}=(p_{o}^{c},T_{o})$ and a trajectory $T_{i}$, the Historical Trajectory Distance $\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$ is no less than $d(p_{o}^{c},p_{i}^{x})$.
Proof.
When computing HTD, it starts with the distance computation between the last points of $T_{o}$ and $T^{\prime}_{i}$, i.e., $p_{o}^{c}$ and $p_{i}^{x}$. Therefore, $d(p_{o}^{c},p_{i}^{x})$ must in the results of max function in Equation 3. If there is an other value in max function results greater than $d(p_{o}^{c},p_{i}^{x})$, the final distance must greater than $d(p_{o}^{c},p_{i}^{x})$. Otherwise, $d(p_{o}^{c},p_{i}^{x})$ will be the final distance of $\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$. ∎
According to Lemma 5.2, an early pruning strategy can be used before fully computing the distance between $o$ and $T_{i}$. $T_{i}$ can be pruned if $d(p_{o}^{c},p_{i}^{x})\geq(\textit{dist}_{k}-(1-\alpha)\cdot\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L}))/\alpha$.
To avoid re-computation, the intermediate information buffer during the search is maintained. It contains a list of information for each computed trajectory $T_{i}$, denoted as $B_{I}=(i,x^{\prime},z^{\prime},\textit{dist}^{\prime}_{H},y^{\prime})$, where $i$ is the ID of $T_{i}$, $x^{\prime}$ is the pivotal point ID in $T_{i}$ at previous timestamp $t^{\prime}$, $z^{\prime}$ is the segment ID of $p_{i}^{x^{\prime}}$ in $T_{i}$, $\textit{dist}^{\prime}_{H}$ is the HTD between $o$ and $T_{i}$ at previous timestamp $t^{\prime}$, and $y^{\prime}$ is the ID of $p_{o}^{y^{\prime}}$ in $T_{o}$, where $\theta^{c-y^{\prime}}*\text{min}\{d(p_{o}^{y^{\prime}},p_{i})|p_{i}\in T^{\prime}_{i}\}=\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$. Then, the search at timestamp $t$ can be accelerated by using the following incremental rule.
Strategy 3 (Incremental Rule).
Given the intermediate buffer of CSTS($o,\mathcal{T},r,k$), if a trajectory $T_{i}$ in $\mathcal{T}$ has been computed in a previous timestamp, the similarity computation at timestamp $t$ can be accelerated by using the incremental rule in different cases instead of fully computing $\mathcal{D}(o,T_{i})$.
Case 1: If the pivotal point ID in $T_{i}$ at $t$ is unchanged, then the TTD $\textit{dist}_{T}$ is unchanged and the HTD $\textit{dist}_{H}$ can be updated as follows.
(9) $\textit{dist}_{H}=\text{max}\{\theta\cdot\textit{dist}^{\prime}_{H},\text{min}\{d(p_{o}^{c},p_{i})|p_{i}\in T^{\prime}_{i}\}\}$
Case 2: If the pivotal point ID in $T_{i}$ at $t$ is greater than $x^{\prime}$, then $\textit{dist}_{T}$ can be computed according to Equation 4. If $\theta^{c-y^{\prime}}\cdot\textit{min}\{d(p_{o}^{y^{\prime}},p_{i})|p_{i}\in\langle p_{i}^{x^{\prime}+1},...,p_{i}^{x}\rangle\}\geq\theta\cdot\textit{dist}^{\prime}_{H}$, where $\langle p_{i}^{x^{\prime}+1},...,p_{i}^{x}\rangle$ is the incremental part of $T^{\prime}_{i}$, then $\textit{dist}_{H}$ can be updated as Equation 9. Otherwise, compute $\textit{dist}_{H}$ according to Equation 3.
Case 3: If the pivotal point ID in $T_{i}$ at $t$ is less than $x^{\prime}$, $\textit{dist}_{H}$, and $\textit{dist}_{T}$ can be computed according to Equations 3 and 4.
Finally, the OTRD can be updated according to Equation 2.
To further speed up the search, we can adjust the granularity of each historical trajectory.
Strategy 4 (Granularity Adjustment).
Given a trajectory $T_{i}$ in $\mathcal{T}$, the pivotal point $p_{i}^{x}$ in $T_{i}$, and a granularity parameter $g$, we have a coarse-grained trajectory $T^{\prime}_{i}(g)=\langle p_{i}^{j},p_{i}^{j+g},\dots,p_{i}^{x-g},p_{i}^{x}\rangle$, where $1\leq j\leq g$. Then, $T^{\prime}_{i}(g)$ will be used in the HTD computation, i.e., $\mathcal{D}_{H}(T_{o},T^{\prime}_{i}(g))$.
However, this strategy may introduce errors to $\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$ since some points in $T_{i}$ are ignored. We deduce the lower and upper error bounds as follows.
(10) $\mathcal{E}^{l}=\mathcal{D}_{H}(T_{o},T^{\prime}_{i})-d(p_{o}^{c},p_{i}^{x})$
(11) $\mathcal{E}^{u}=\text{max}\{\theta^{c-j}\cdot d(p_{o},p_{i}^{j})|p_{o}\in T_{o},p_{i}^{j}\in T^{\prime}_{i}(g)\}-\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$
The lower error bound $\mathcal{E}^{l}$ can be easily proved according to Lemma 5.2. The upper error bound can be proved as follows.
Proof.
If $\exists p_{i}^{j}\in T^{\prime}_{i}\setminus T^{\prime}_{i}(g)$, $\exists p_{o}\in T_{o}$, such that $\theta^{c-j}\cdot d(p_{o},p_{i}^{j})>\textit{max}\{\theta^{c-j}\cdot d(p_{o},p_{i}^{j})|p_{o}\in T_{o},p_{i}^{j}\in T^{\prime}_{i}(g)\}$, then $d(p_{o},p_{i}^{j})$ must be excluded in min function in Equation 3. Therefore, the upper bound of $\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$ is $\textit{max}\{\theta^{c-j}\cdot d(p_{o},p_{i}^{j})|p_{o}\in T_{o},p_{i}^{j}\in T^{\prime}_{i}(g)\}$. ∎
To apply the incremental rule to $T^{\prime}_{i}(g)$, we design the update rule of $T^{\prime}_{i}(g)$ at timestamp $t$ with the intermediate buffer $B_{I}$ when the pivotal point is changed from $p_{i}^{x^{\prime}}$ to $p_{i}^{x}$.
[…]
(13) $\gamma=|x-x^{\prime}|\mbox{ mod }g$
The proposed speed-up strategies can enhance the efficiency of continuous similar trajectory search by leveraging mathematical properties of trajectory segments and intermediate results. Segment pruning and $k$-bound pruning effectively eliminate irrelevant candidates early in the process, while the incremental rule optimizes computations through reuse of prior results. Granularity adjustment further reduces computational overhead by coarsening trajectory representation within controlled error bounds. Together, these strategies provide a robust framework to accelerate similarity computations, enabling real-time and scalable trajectory analysis.
5.2. OTRD Computation
The similarity computation between a moving object $o$ and a trajectory $T_{i}$ is performed in several key steps. First, the algorithm identifies the pivotal point $p_{i}^{x}$ of $T_{i}$ to the current position $p_{o}^{c}$ of $o$. Next, it computes the distance between $o$’ target destination $p_{o}^{d}$ and the line $L(p_{i}^{x},p_{i}^{n})$, where $p_{i}^{n}$ is the end point of $T_{i}$. Afterward, the distance between the moving object’s current trajectory $T_{o}$ and the sub-trajectory $T^{\prime}_{i}=\langle p_{i}^{1},...,p_{i}^{x}\rangle$ from $T_{i}$. To enhance the efficiency during this process, the speed-up strategies are employed to avoid unnecessary distance computations. Finally, the overall similarity is determined by summing these distances with a trade-off parameter $\alpha$, providing a comprehensive measure of how closely the moving object’s trajectory aligns with the historical one.
Algorithm 1 OTRD_Computation($o$, $T_{i}$, $\alpha$, $\theta$, $g$, $\textit{dist}_{k}$, $B_{I}$)
1:Input: A moving object $o$, a trajectory $T_{i}$, the trade-off parameter $\alpha$, the decay parameter $\theta$, the granularity parameter $g$, the current top-$k$ threshold $\textit{dist}_{k}$, and the intermediate buffer $B_{I}$.
2:Output: The distance between $o$ and $T_{i}$, and the updated intermediate buffer $B_{I}$.
[…]
6:if $B_{I}[T_{i}].x^{\prime}$ ¡ $x$ then
7:  $\textit{dist}_{n}\leftarrow$ Min_Dist($B_{I}[T_{i}].y^{\prime}$, $B_{I}[T_{i}].z^{\prime}$, $z$, $T$, $B_{I}[T_{i}].\textit{dist}^{\prime}_{H}$, $g$)
8:  if $\textit{dist}_{n}>B_{I}[T_{i}].\textit{dist}^{\prime}_{H}$ then $\textit{isValid}\leftarrow\textit{true}$
[…]
10:for each point $p_{o}^{j}$ in $T_{o}$ do
11:  if $p_{o}^{j}\neq p_{o}^{c}$ and isValidisValid then continue $\triangleright$ Strategy 3
12:  $\textit{dist}_{n}\leftarrow$ Min_Dist($j$, 0, $z$, $T$, $\infty$, $g$)
13:  if $\textit{dist}_{H}<\theta^{c-j}\cdot\textit{dist}_{n}$ then
14:    $\textit{dist}_{H}\leftarrow\theta^{c-j}\cdot\textit{dist}_{n}$, $p_{\textit{max}}\leftarrow j$
15:    $\textit{dist}^{\prime}_{k}\leftarrow(\textit{dist}_{k}-(1-\alpha)\cdot\textit{dist}_{T})/(\alpha\cdot\theta^{c-j})$
16:    if $\textit{dist}_{H}>\textit{dist}^{\prime}_{k}$ then break $\triangleright$ Strategy 2
[…]
19:return dist and $B_{I}$
20:function Min_Dist($j$, $^{\prime}, zz$, $T_{i}$, $\textit{dist}_{n}$, $g$)
21:  for each segment $s_{i}^{u}$ in $T_{i}$ with $u\in(z^{\prime},z]$ do
22:    $\textit{dist}_{s}\leftarrow\textsc{Compute\_Dist\_S}(p_{o}^{j},s_{i}^{u})$
23:    if $\textit{dist}_{s}>\textit{dist}_{n}$ then continue $\triangleright$ Strategy 1
24:    for each point $p_{i}$ in $s_{i}^{u}$ with $g$ do $\triangleright$ Strategy 4
25:     $\textit{dist}_{p}\leftarrow d(p_{o}^{j},p_{i})$, $\textit{dist}_{n}\leftarrow\textit{min}(\textit{dist}_{n},\textit{dist}_{p})$
26:  return $\textit{dist}_{n}$
Algorithm 1 shows the details of the similarity computation. Specifically, we first identify the nearest segment $s_{i}^{z}$ and the nearest point $p_{i}^{x}$ on the historical trajectory $T_{i}$ to the moving object’s current position $p_{o}^{c}$ (lines 1–2). We initialize a boolean flag isValid to check if the pre-computed distance in the intermediate buffer $B_{I}$ is still valid (line 3). A variable $p_{\textit{max}}$ is also initialized to represent the ID of the point $p_{o}^{y}$ in $T_{o}$, where $\theta^{c-y}*\textit{min}\{d(p_{o}^{y},p_{i})|p_{i}\in T^{\prime}_{i}\}=\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$ (line 3). If the pivotal point ID of $T_{i}$ increases and the distance between $p_{o}^{y^{\prime}}$ and the incremental part is greater than the maintained distance, isValid is set to true (lines 4–6). Then, the TTD $\textit{dist}_{T}$ is computed, which is the distance between the moving object’s target destination $p_{o}^{d}$ and a line segment on the historical trajectory. (line 7) Additionally, the HTD $\textit{dist}_{H}$ is initialized as 0 (line 7). Then, $\textit{dist}_{H}$ is computed (lines 8–14). If the maintained distance is valid, we only update the distance between $p_{o}^{c}$ and $T^{\prime}_{i}$ (Case 2 in Strategy3). Otherwise, $\textit{dist}_{H}$ should be computed by traversing every point $p_{o}^{j}$ in $T_{o}$ (from $p_{o}^{c}$ to previous points) to compute the minimum distance between $p_{o}^{j}$ and $T^{\prime}_{i}$ using Min_Dist function, where the segment pruning (Strategy 1) and the granularity adjustment (Strategy 4) are used during the computation(lines 18–24). During the traverse, the $p_{\textit{max}}$ is updated (line 12) and the $k$-bound pruning (Strategy 2) is used to eliminate irrelevant candidates early (lines 13–14). Then, the OTRD value is computed by summing $\textit{dist}_{H}$ and $\textit{dist}_{T}$ with a trade-off parameter $\alpha$ (line 15). Next, the intermediate buffer is updated (line 16). Finally, the OTRD value and the updated intermediate buffer are returned (line 17).
5.3. CSTS Algorithm
The CSTS algorithm leverages the index SVTI, the similarity measure OTRD, and four speed-up strategies to efficiently identify the top $k$ similar trajectories for a moving object $o$ at each timestamp $t$.
Algorithm 2 outlines the CSTS algorithm. Initially, two parameters are set: the decay parameter $\theta$, which controls the influence of past trajectory points on similarity computations, and the trade-off parameter $\alpha$, which balances the weights of the two components in OTRD (lines 1–2). Next, for each timestamp $t$, we first initialize an intermediate buffer $B_{I}$ to maintain intermediate results (Strategy 3) and a min-heap result buffer $B_{k}$ to keep the computed similarity values for each historical trajectory at timestamp $t$ (lines 4–5). Afterwards, a range query is executed to find all candidates $\mathcal{T}_{r}$ using the index SVTI (line 6). For each candidate trajectory $T_{i}$, it computes the similarity between $T_{i}$ and $o$ (lines 8–9). The current $k$-th smallest similarity value $\textit{dist}_{k}$ is retrieved from $B_{k}$. The similarity is computed using the OTRD_Computation function incorporating pruning strategies such as segment pruning and $k$-bound pruning. The resulting similarity value and trajectory ID are inserted into $B_{k}$ (line 10). After evaluating all candidate, the top $k$ trajectories with the smallest OTRD values are extracted from $B_{k}$ (line 11). Finally, the top $k$ results for the current timestamp $t$ are returned (line 12).
Algorithm 2 CSTS_Algorithm($o$, $\mathcal{T}$, $r$, $k$, $g$, SVTISVTI)
1:Input: A moving object $o$, a historical trajectory set $\mathcal{T}$, a range $r$, an integer $k$, the granularity parameter $g$, segment-based vessel trajectory index SVTISVTI.
2:Output: Top $k$ similar trajectories in $\mathcal{T}$ at each timestamp.
3:Initialize a decay parameter $\theta$
4:Initialize a trade-off parameter $\alpha$
5:for each timestamp $t$ do
6:  Initialize an intermediate buffer: $B_{I}\leftarrow\emptyset$
7:  Initialize a min-heap result buffer: $B_{k}\leftarrow\emptyset$
8:  $\mathcal{T}_{r}\leftarrow\textsc{Range}(o,\mathcal{T},r,k,SVTI)$
9:  for each candidate trajectory $T_{i}$ in $\mathcal{T}_{r}$ do
10:    $\textit{dist}_{k}\leftarrow B_{k}.\textit{pop}()$
11:    $\textit{dist}\leftarrow$ OTRD_Computation($o$, $T_{i}$, $\alpha$, $\theta$, $g$, $\textit{dist}_{k}$, $B_{I}$)
12:    Insert (dist, $T_{i}$) into $B_{k}$
13:  $\Theta\leftarrow\textsc{Get\_K}(B_{k})$
14:  return $\Theta$ at timestamp $t$
6. Experiments
6.1. Overall Experimental Settings
All experiments were conducted on a Linux server running Ubuntu with an Intel Xeon Icelake processor featuring 60 cores at 5.7 GHz and 240 GB of memory. The system is equipped with dual NVIDIA A10 GPUs, each providing 23 GB VRAM, for a total of 46 GB of GPU memory. All neural network models are implemented using PyTorch and trained on the GPU, and all index construction and search algorithms are implemented in Python (Code for CSTS [^1] https://github.com/hyLiu1994/CSTS/blob/main/README.md (Last access: 2025.1)).
Datasets. We evaluate the proposed methods on the following datasets. Dataset statistics are reported in Table 2.
Tasks. We conduct extensive experiments to address and analyze the following research questions (RQs):
  • RQ1. Evaluation on SVTI: How efficient is the construction process of SVTI?
  • RQ2. Evaluation on CSTS: How does the search algorithm CSTS perform in terms of efficiency and effectiveness?
  • RQ3. Ablation Study: What is the impact of the speed-up strategies on overall performance?
Table 2. Dataset statistics
  Dataset Size (GB) Trajectories Points
DK-AIS 37.43 746,302 216,189,595
US-AIS 15.61 737,133 154,279,662
Table 3. Metrics and baselines
  Questions Metrics Baselines
RQ1 Construction Time (s), Index Size (MB) O-PQT, O-PRT, H/F-TrajI, H/F-DFT
RQ2 Query Time (ms), Hit (%) O-PQT, O-PRT, H/F-TrajI, H/F-DFT, TrajCL-TI, CLEAR-TI
[…]
Metrics. The metrics for each task are listed in Table 3. To evaluate the efficiency of SVTI, we measure its construction time and index size. To evaluate the performance of CSTS, we select ten query instances, execute the selected queries ten times, and report the average. For effectiveness evaluation, we employ the hit rate ($\mathit{Hit}$):
(14) $\mathit{Hit}=\frac{\lvert\mathcal{T}_{f}\cap\mathcal{T}\rvert}{k},$
where $\mathcal{T}_{f}$ is the set of trajectories that cover the $k$-nearest neighbor (kkNN) points of moving object $o$’s next future point and $\mathcal{T}$ is the query result.
Baselines. We compare the proposed methods with several baseline methods, as outlined in Table 3. Specifically, we study SVTI and four traditional techniques with respect to index construction time and size. Furthermore, CSTS is compared with six traditional methods and two learning-based methods to evaluate its efficiency and effectiveness. The baseline methods are listed as follows:
  • O-PQT: OTRD + PQT. A continuous similar trajectory search method on a Point-based Quad-Tree (PQT) using the OTRD measure.
  • O-PRT: OTRD + PRT. A continuous similar trajectory search method on a Point-based R-Tree (PRT) using the OTRD measure.
  • H-TrajI/F-TrajI: Hausdorff/Fréchet + Traj_Index (Xie et al., 2017). An R-Tree-based similar trajectory method using Hausdorff or Fréchet distance, where the index is built according to the centroid of each data trajectory’s minimum bounding rectangle.
  • H-DFT/F-DFT: Hausdorff/Fréchet + DFT (Xie et al., 2017). A Distributed Framework for Trajectory similarity search (DFT) using Hausdorff or Fréchet distance.
  • TrajCL-TI: TrajCL (Chang et al., 2023) + SVTI. A trajectory similarity search method based on SVTI, employing a graph-based long-term dependency model for similarity computation.
  • CLEAR-TI: CLEAR (Li et al., 2024a) + SVTI. A trajectory similarity search method using SVTI with a similarity computation technique based on contrastive representation learning.
[…]
Parameters. The parameter settings are summarized in Table 4, with default values in bold. A description of each parameter is provided below:
  • $l_{\textit{min}}$ and $l_{\textit{max}}$: The minimum and maximum segment lengths.
  • Data size ($\Lambda$): The number of the points in the AIS data. Five different data sizes are evaluated: D-1, D-2, D-3, D-4, and D-5, corresponding to approximately $20\%$, $40\%$, $60\%$, $80\%$, and $100\%$ of the original dataset.
  • Query length ($l_{q}$): The length of the historical trajectory of moving object $o$ at the query’s starting time $t$.
  • Range ($r$): The range of the moving object $o$’s current position at each timestamp $t$. This is represented by the candidate rate, e.g., a range of 10 indicates that the candidate number is approximately $10\cdot k$.
  • Top $k$: The number of similar trajectories returned.
  • Decay factor ($\theta$): The decay factor used in the object-trajectory real-time distance computation.
  • Trade-off factor ($\alpha$): The trade-off factor in the object-trajectory real-time distance computation.
  • Granularity ($g$): The granularity used in Strategy 4.
For each query, we perform a continuous similar trajectory search across 20 timestamps and report the average results.
Table 4. Parameter settings
  Parameters Questions Settings
$l_{\textit{min}}$ RQ1 20, 25, 30, 35, 40
$l_{\textit{max}}$ RQ1 40, 45, 50, 55, 60
$\Lambda$ RQ1, RQ2 D-1, D-2, D-3, D-4, D-5
$l_{q}$ RQ2 10, 20, 30, 40, 50
$r$ RQ2 5, 10, 15, 20, 25
$k$ RQ2 30, 40, 50, 60, 70
$\theta$ RQ2 0.45, 0.5, 0.55, 0.6, 0.65
$\alpha$ RQ2 0.45, 0.5, 0.55, 0.6, 0.65
$g$ RQ3 1, 5, 10, 15, 20
6.2. Evaluation on CSTS
We investigate the query time and hit rate of CSTS, and the baselines under different parameter settings.
(a) Time vs. Dataset
(b) Hit vs. Dataset
Figure 6. Comparison of different methods on querying
6.2.1. Default
Figure 6(a) presents the query time for all methods on the two datasets under the default settings. CSTS demonstrates the best performance in terms of query time, attributed to its well-designed search algorithm with speed-up strategies. O-PRT has the longest search time, approximately 57 seconds on DK-AIS and 13 seconds on US-AIS, and its results are not displayed in the figure due to the significant disparity. H-TrajI and F-TrajI, which utilize an index based on entire trajectories, incur higher costs during the candidate search phase within the specified range. H-DFT and F-DFT perform relatively better than H/F-TrajI because the segment-based index offers increased efficiency during candidate searching. Both H/F-TrajI and H/F-DFT employ the Hausdorff or Fréchet distance measures, which increase the search time during distance computations considerably, as they consider all point pairs within the trajectories. TrajCL-TI and CLEAR-TI, the two AI-based methods, exhibit strong efficiency since they leverage GPU capabilities, enabling faster distance computations. O-PQT takes longer than CSTS, indicating that the SVTI in CSTS is more effective than the Quad-tree based index.
Figure 6(b) shows the hit rates of all methods on the two datasets under default settings. The hit rate of CSTS reaches 81% on DK-AIS and 87% on US-AIS, the highest among all methods. This superior performance is attributed to the proposed OTRD measure, which incorporates future movement trends of trajectories. O-PQT and O-PRT exhibit hit rates identical to that of CSTS, as they also employ the OTRD measure, which determines effectiveness in similar trajectory search. H-TrajI and H-DFT achieve the same hit rate because both use Hausdorff distance, while F-TrajI and F-DFT also perform equivalently in terms of hit rate due to their use of Fréchet distance. The Hausdorff and Fréchet measures show comparable hit rates but perform worse than OTRD because they rely on static or full-path comparisons, making them a poor fit for evolving trajectories. The two AI-based similarity computation methods, TrajCL-TI and CLEAR-TI, achieve the lowest hit rates because they are designed to cluster trajectory variants and are not optimized for computing similarity across all original trajectories.
(a) Time vs. $l_{q}$ (DK-AIS)
(b) Hit vs. $l_{q}$ (DK-AIS)
(c) Time vs. $l_{q}$ (US-AIS)
(d) Hit vs. $l_{q}$ (US-AIS)
Figure 7. Effect of $l_{q}$
6.2.2. Effect of $l_{q}$
As shown in Figures 7(a) and 7(c), the query times for all methods increase as the query length $l_{q}$ increases. Among the traditional methods, CSTS exhibits the smallest rate of increase because it only evaluates the historical portion of trajectories, avoiding point-by-point checks for the entire trajectory. The query time for O-PQT also increases slightly since it uses the same similarity measure as CSTS, but its reliance on a point-based quad-tree yields higher search costs. H/F-TrajI and H/F-DFT exhibit a more pronounced growth in query time as $l_{q}$ increases, as each point in the historical part of the query trajectory is compared against the entire set of historical trajectories. In contrast, TrajCL-TI and CLEAR-TI show almost no variation in query time varying $l_{q}$, as these methods transform trajectories into fixed-length vectors, making them independent of query length. Figures 7(b) and 7(d) reveal no discernible trend in the hit rate, suggesting that the effect of $l_{q}$ on hit rate is relatively minor.
6.2.3. Effect of $r$
As shown in Figures 8(a) and 8(c), the query time for all methods increases with the range $r$, as a larger range results in more candidates to evaluate. H/F-TrajI and H/F-DFT experience the fastest growth in query time due to the higher computational cost of similarity computations for each candidate compared to OTRD and the AI-based methods. Among all methods, CSTS remains the most efficient, even as the range increases, owing to its optimized search algorithm. Figures 8(b) and 8(d) show that the hit rate decreases as the range increases. This is because a larger range includes more candidates. Some of these candidates might appear in the final similar trajectory results but not in the kkNN of $o$’s next position, leading to a reduced hit rate.
(a) Time vs. $r$ (DK-AIS)
(b) Hit vs. $r$ (DK-AIS)
(c) Time vs. $r$ (US-AIS)
(d) Hit vs. $r$ (US-AIS)
Figure 8. Effect of $r$
(a) Time vs. $k$ (DK-AIS)
(b) Hit vs. $k$ (DK-AIS)
(c) Time vs. $k$ (US-AIS)
(d) Hit vs. $k$ (US-AIS)
Figure 9. Effect of $k$
6.2.4. Effect of $k$
As shown in Figure 9, increasing the value of $k$ leads to trends in query time and hit rate similar to those observed when increasing the value of $r$. Specifically, the query time increases, while the hit rate decreases. This occurs because the number of candidates is directly proportional to $k$, and a larger $k$ results in more candidates to evaluate, thereby increasing the time required for similarity computation. Moreover, the hit rate decreases as $k$ grows since $k$ serves as the denominator in the hit rate computation, making it more likely to produce a lower hit rate with larger $k$.
6.2.5. Effect of Data Size
We evaluate CSTS under varying data sizes, as shown in Figure 10. The settings for datasets D-1 through D-5 remain consistent with those described in Section B.1.2. As the data size increases, both the query time and the hit rate remain relatively stable. This is because the number of candidates remains consistent regardless of the dataset size. Consequently, the query time is primarily determined by the similarity computation rather than the candidate-finding phase, which incurs negligible overhead.
(a) Time vs. $\Lambda$ (US-AIS)
(b) Hit vs. $\Lambda$ (US-AIS)
Figure 10. Effect of data size
6.2.6. Effect of $\theta$ and $\alpha$
We examine the impact of the decay factor $\theta$ and the trade-off factor $\alpha$ on CSTS. Since these parameters do not influence efficiency, only their effects on the hit rate are reported. As shown in Figure 11(a), the hit rate decreases slightly as $\theta$ increases for both datasets. This is because a smaller $\theta$ results in a higher decay for earlier points in the historical portion of the query object. Consequently, points closer to the current position of the query object are given higher weight during similarity computation, leading to more accurate trajectory matches.
Similarly, Figure 11(b) shows a slight decline in the hit rate as $\alpha$ increases. This indicates that the Historical Trajectory Distance plays a more significant role in the similarity measure compared to the coarser-grained Target-Trajectory Distance.
(a) Hit vs. $\theta$
(b) Hit vs. $\alpha$
Figure 11. Effect of $\theta$ and $\alpha$
7. Related Work
Trajectory similarity measures. Trajectory similarity measures form the foundation of trajectory analysis. Dynamic Time Warping (Yi et al., 1998; Ying et al., 2016) allows for flexible alignment and comparison of time-series data, accommodating variations in speed and temporal shifts. Longest Common Subsequence (Vlachos et al., 2002) offers robustness to noise and spatiotemporal deformations by identifying the longest subsequence shared between trajectories. Additionally, Fréchet distance (Alt and Godau, 1995; De Berg et al., 2013; Agarwal et al., 2014) and edit distance-based measures (Chen and Ng, 2004; Chen et al., 2005; Ranu et al., 2015) provide nuanced assessments of trajectory similarity, considering geometric constraints and structural differences. The Hausdorff distance (Hangouet, 1995; Bai et al., 2011) measures how far the points of one set are from the points of the other set. These measures (Toohey and Duckham, 2015; Magdy et al., 2015) serve as the cornerstone for trajectory analysis. However, these measures fail to consider the significance of individual points, the trends within historical trajectories, and the future movement patterns of objects.
Learning-based trajectory similarity computation. Recent advancements in trajectory similarity search have explored learning-based approaches (Li et al., 2018; Fu and Lee, 2020; Liu et al., 2020; Deng et al., 2022; Fang et al., 2022; Mao et al., 2022; Li et al., 2024a) to improve the accuracy and efficiency of similarity computations. These methods often leverage deep learning models to automatically learn representations of trajectories from raw spatiotemporal data. However, learning-based trajectory similarity computation has several limitations that make it less suitable for real-time, continuous vessel trajectory search. First, these methods often rely on down-sampling or distortion of trajectories, which leads to the identification of trajectories as similar, even when they lack true alignment with the original paths. This compromises their ability to accurately capture the underlying similarity between the original trajectories. Second, these methods typically focus on computing the similarity between two given trajectories rather than provide efficiently similarity search method on a large trajectory database. Third, they are not well-suited for real-time processing. The continuous updates in vessel trajectories would necessitate frequent model re-training or recalibration, imposing substantial computational overhead and significantly degrading performance. Finally, learning-based models are generally designed for offline batch processing, making them unsuitable for dynamic and real-time maritime navigation.
Similar trajectory search. Studies on similar trajectory search (Ding et al., 2008; Xie et al., 2017; Wang et al., 2018; Li et al., 2020; He et al., 2022; Luo and Zeng, 2023) and similarity join (Bakalov et al., 2005b, a; Ta et al., 2017; Shang et al., 2017) have proposed various techniques to identify similar movement patterns. Some approaches (Hwang et al., 2006, 2005; Xie et al., 2017; Yuan and Li, 2019; Han et al., 2021; He et al., 2022; Fang et al., 2022) are widely used in domains like vehicle tracking and urban traffic monitoring, particularly in road networks, where vehicle movements are constrained by structured paths, making it easier to organize and index trajectories based on fixed road segments, but not suitable for vessel trajectories. Furthermore, they mainly designed for existing trajectory similarity methods to compare entire trajectories, which is not suitable for real-time similar vessel trajectory search. These methods are less effective for real-time applications, where the trajectory evolves over time and must be analyzed incrementally. As a result, current trajectory similarity search methods are not well-suited to solve the problem of real-time, continuous vessel trajectory search.
8. Conclusion and Future Work
This paper presents ACTIVE, a robust framework for real-time continuous trajectory similarity search for vessels. By introducing the OTRD measure, ACTIVE focuses on predictive, forward-looking trajectory analysis, overcoming the limitations of traditional retrospective methods. We propose the CSTS algorithm for continuous search of similar historical trajectories. The developed index, SVTI, along with four speed-up strategies, enhances the system’s efficiency and scalability, enabling rapid and accurate similarity computations. Experiments on real-world AIS datasets demonstrate that ACTIVE outperforms state-of-the-art methods by up to 70% in terms of query time, 60% in terms of hit rate, and reduces index construction costs significantly.
In future, it is of interest to apply the proposed continuous similarity search to real-time maritime applications, such as traffic monitoring, collision avoidance, and route optimization, enabling dynamic and proactive decision-making in large-scale vessel trajectory analysis.
Acknowledgements.
We acknowledge the support of the EU’s Horizon Europe research program under grant agreement No. 101070279 (MobiSpaces).
Appendix A Algorithm of Segment Partitioning
Algorithm 3 outlines the segment partitioning process in detail. Initially, a segment set $\mathcal{S}$ is initialized (line 1). For each trajectory $T_{i}$, the algorithm identifies the optimal segmentation (lines 2–16). Specifically, a list is created for $T_{i}$’s segment set, where the $j$th item represents the current optimal segment set up to point $p^{j}$ (line 3). Additionally, another list is initialized to store the total MBR area, where the $j$th item represents the minimum total MBR area up to point $p^{j}$ (line 4). For each point $p^{j}$ in $T_{i}$, the minimum total MBR area up to $p^{j}$ is computed (lines 5–15). If $j$ is less than $l_{\textit{min}}$, the area is set to 0; otherwise, it is initialized to $\infty$ (lines 6–7). For each possible starting point $p^{j-l+1}$, where $l_{\textit{min}}\leq l\leq l_{\textit{max}}$, the MBR area of the segment $\langle p^{j-l+1},...,p^{j}\rangle$ is computed (lines 8–10). If the total MBR area is smaller than the current value, both the area and the segment set are updated (lines 11–15). Once all points in $T_{i}$ have been processed, the final segment set for $T_{i}$ is added to the overall segment set $\mathcal{S}$ (line 16). Ultimately, the algorithm returns the overall segment set $\mathcal{S}$ (line 17).
Algorithm 3 Segment_partitioning($\mathcal{T}$, $l_{\textit{max}}$, $l_{\textit{min}}$)
1:Input: A vessel trajectory set $\mathcal{T}$, the maximum segment length $l_{\textit{max}}$, and the minimum segment length $l_{\textit{min}}$.
2:Output: A segment set $\mathcal{S}$.
3:Initialize a segment set $\mathcal{S}$ $\leftarrow$ $\emptyset$
4:for $T_{i}=\langle p_{i}^{1},p_{i}^{2},...,p_{i}^{n}\rangle$ in $\mathcal{T}$ do
5:  Initialize a list for $T_{i}$’s segment set $S_{i}$ $\leftarrow$ $\emptyset$
6:  Initialize a list for the total MBR area $F$ $\leftarrow$ $\emptyset$
7:  for $j$ in $[1,n]$ do
8:   if $j\leq l_{\textit{min}}$ then $F[j]$ $\leftarrow$ 0
9:   else $F[j]$ $\leftarrow$ $\infty$
10:   for $l$ in $l_{\textit{min}},l_{\textit{max}}]$ do
11:     if $j-l\geq 0$ then
12:      $m_{l}$ $\leftarrow$ $\mathcal{M}_{area}$($p_{i}^{j-l+1}$, $p_{i}^{j}$)
13:      if jlj-l is 0 then $m$ $\leftarrow$ $m_{l}$
14:      else $m$ $\leftarrow$ $F[j-l]+m_{l}$
15:      if $F[j]>m$ then
16:        $F[j]$ $\leftarrow$ $m$
17:        $S_{i}[j]$ $\leftarrow$ $S_{i}[j-l]$ $\cup$ $\langle p_{i}^{j-l+1},...,p_{i}^{j}\rangle$
18:  $\mathcal{S}\leftarrow\mathcal{S}\cup S_{i}[n]$
19:return $\mathcal{S}$
Appendix B Evaluation
B.1. Evaluation on SVTI
B.1.1. Default
The performance of the different index methods under the default parameter settings is presented in Figure 12. In terms of construction time, SVTI takes 42 minutes to build on the DK-AIS dataset and 40 minutes on the US-AIS dataset. The index sizes of SVTI are 196 MB and 235 MB on DK-AIS and US-AIS, respectively. Among the five indices, PRT has the longest construction time, exceeding one day for both datasets due to its point-based design. Consequently, PRT also has the largest index size. Next, TrajI is the most efficient index to construct, as it is based on entire trajectories. This design minimizes insertion times, making its construction time directly proportional to the number of insertions. The original segmentation method for DFT, designed for road networks, cannot be applied directly to AIS data, so our segment partitioning method is used. Therefore, the construction time and index size of DFT are comparable to those of SVTI. Lastly, PQT exhibits relatively higher efficiency compared to PRT because the Quad-tree based PQT benefits from faster theoretical construction times compared to the R-tree based PRT.
(a) Time vs. Dataset
(b) Size vs. Dataset
Figure 12. Comparison of index construction time and size
B.1.2. Effect of Data Size.
We analyze the effect of varying data sizes on construction time and index size, with results shown in Figures 13(a) and 13(b), respectively. The datasets D-1, D-2, D-3, D-4, and D-5 contain approximately 43, 86, 130, 172, and 216 million points for DK-AIS, and 31, 62, 93, 123, and 154 million points for US-AIS. As the data size increases, the number of segments and corresponding insertions also grow. This leads to a proportional increase in both construction time and index size.
(a) Time vs. $\Lambda$
(b) Size vs. $\Lambda$
Figure 13. Effect of data size
B.1.3. Effect of $l_{\textit{max}}$ and $l_{\text{min}}$
(a) Time vs. $l_{\text{min}}$
(b) Size vs. $l_{\text{min}}$
Figure 14. Effect of $l_{\text{min}}$ ($l_{\textit{max}}=l_{\text{min}}+20$)
(a) Time vs. $l_{\textit{max}}$
(b) Size vs. $l_{\textit{max}}$
Figure 15. Effect of $l_{\textit{max}}$ ($l_{\text{min}}=30$)
We evaluate the efficiency of SVTI construction when varying different segmentation parameter settings. First, we vary the minimum segment length $l_{\text{min}}$, keeping the maximum segment length $l_{\textit{max}}$ at $l_{\text{min}}+20$. As $l_{\text{min}}$ increases, the time consumption in the segmentation phase, i.e., DK-AIS-S and US-AIS-S, shows a slight increase for across both datasets. This occurs because larger values of $l_{\text{min}}$ require more iterations in the segmentation process. However, the total construction time and index size decrease as $l_{\text{min}}$ increases. This is due to the generation of fewer segments, which subsequently reduces both the time required for construction and the overall index size.
We also vary the maximum segment length $l_{\textit{max}}$, while fixing the minimum segment length $l_{\text{min}}$ at 30. The results are shown in Figure 15. An increase in $l_{\textit{max}}$ results in longer segmentation times for both datasets, as the process requires more iterations with larger $l_{\textit{max}}$ values. Additionally, the number of generated segments decreases slightly as $l_{\textit{max}}$ grows, leading to a modest reduction in both index size and the construction time during the building phase, i.e., DK-AIS-B and US-AIS-B.
B.2. Ablation Study
By default, CSTS incorporates segment pruning (Strategy 1), $k$-bound pruning (Strategy 2), and the incremental rule (Strategy 3). These three strategies are solely designed to improve efficiency and do not reduce the hit rate. In Figure 16, we report their impact on query time. Among the strategies, the incremental rule (S3) proves to be the most effective, reducing the query time by at least threefold. This is the evidence of the efficiency of the incremental rule in optimizing CSTS. The second-most effective strategy is $k$-bound pruning (S2), which shows comparable performance to S3. Segment pruning (S1), however, provides relatively less improvement on both datasets. This is because it introduces additional computational costs to compute the distance between a point and the minimum bounding rectangle of a segment.
The granularity adjustment (Strategy 4) is not applied in CSTS by default, as it may introduce errors. To evaluate the effectiveness of this strategy, we vary the granularity parameter $g$. Referring to Figure 17, both query time and hit rate decrease as $g$ increases. With a larger $g$, more points are disregarded during similarity computation, resulting in reduced query time. However, the hit rate declines because ignoring points introduces errors.
B.3. Case Study on Similarity Measures
Figure 18 presents case studies conducted on DK-AIS and US-AIS, illustrating the differences in results when applying different similarity measures. The orange line represents a moving object $o$, including both its current trajectory and the projected destination. The black line indicates the most similar trajectory identified using the OTRD measure. In comparison with other similarity measures such as Hausdorff, Fréchet, TrajCL, and CLEAR, the OTRD measure demonstrates a clear advantage. The other four measures focus predominantly on aligning with $o$’s current trajectory, resulting in matches that are less aligned with $o$’s overall path. Conversely, the trajectory identified by OTRD not only aligns closely with $o$’s current trajectory but also incorporates the future trend and destination, offering a more holistic similarity result. This highlights the strength of OTRD at capturing the dynamic and predictive aspects of trajectory comparison, making it suitable for applications where future trends and endpoints are important.
Figure 16. Effect of speed-up strategies
(a) Time vs. $g$
(b) Hit vs. $g$
Figure 17. Effect of granularity
Figure 18. Case study